Solve the following by substitution: x + y = 0, 3x + 3y = 4. Are these equations independent, dependent or inconsistent?
What are the slopes of the above lines ?
Change the second equation in problem 1 to 3x + 2y = 4. Solve by substitution.

Answer by Alan3354(69443)

Solve the following by substitution: x + y = 0, 3x + 3y = 4. Are these equations independent, dependent or inconsistent?
------------------
They're inconsistent. Multiply the 1st eqn by 3:
--> 3x + 3y = 0
The 2nd says 3x + 3y = 4
If you graph them, they're parallel. No intersection, no solution.
--------------------------
What are the slopes of the above lines ?
The slopes are equal, = -1
----------------------------
Change the second equation in problem 1 to 3x + 2y = 4. Solve by substitution.
x + y = 0 --> y = -x
3x + 2y = 4
Sub for y
3x + 2(-x) = 4
x = 4
y = -4